OpenStudy (anonymous):

Simplify completely:

OpenStudy (anonymous):

\[10x^6y^3+20x^3y^2 \over 5x^3y\]

OpenStudy (anonymous):

Because your numerators aren't like terms, I'd start by giving each numerator its own denominator, like this:\[\frac{10x^6y^2}{5x^3y} + \frac{20x^3y^2}{5x^3y}\]Then, simplify each fraction individually.

OpenStudy (anonymous):

That would leave you with this: \[\frac{2x^3y}{1} + \frac{4y}{1}\] Which is simply:\[2x^3y+4y\]

OpenStudy (anonymous):

thanks a lot!

OpenStudy (anonymous):

Happy to help!